1.     Generate 4 random numbers by using linear congruenital generator.Use X0 = 21, a= 17, c=43, m=105.


1
Expert's answer
2021-06-01T14:51:51-0400

"X_0=21."

"X_n=(aX_{n-1}+c)\;mod \;m".

"X_1=(17*21+43)\;mod\;105=400\;mod\;105=85."

"X_2=(17*85+43)\;mod\;105=1488\;mod\;105=18."

"X_3=(17*18+43)\;mod\;105=349\;mod\;105=34."

"X_4=(17*34+43)\;mod\;105=621\;mod\;105=96."